# K theoretic virtual fundamental cycle of global Kuranishi chart

### v1

@@ Abstract.
In this paper, we define the virtual fundamental cycle of a global Kuranishi chart, as an element in the (analytic) orbifold K-homology of the base space of the chart, and verify that it defines the same invariants as in [2].
@@ 1. Introduction
The motivation of this paper is to define permutation equivariant quantum K-theory [12][6][7] for general Symplectic target spaces. The essential ingredient in the definition is the $K$-theoretical virtual fundamental cycle of moduli spaces $\mathcal{M}$ of J-holomorphic curves. These moduli spaces are characterized using almost complex global Kuranishi charts [1][2].
@@
In this paper, we define $K$-theoretical virtual fundamental cycles as elements in the $K$-homologies of certain spaces, instead of $K$-cohomologies of cotangent spaces as in [1][2].
@@ 1.1. The case of manifolds
Let $Y$ be a compact almost complex manifold (possibly with boundary), and $s$ a section of a complex vector bundle $E$ over $Y$ whose zero locus $s^{-1}(0)$ lies in the interior of $Y$. Pick Hermitian metrics on $T^{\ast}Y$ and $E$.
Let $\pi:T^{\ast}Y\to Y$ and
$\begin{array}[]{llll}S_{+}&=\wedge^{even}T^{\ast}Y,&S_{-}&=\wedge^{odd}T^{\ast}Y,\\ E_{+}&=\wedge^{even}E,&E_{-}&=\wedge^{odd}E.\end{array}$
@@
Pick a vector bundle $O^{\prime}$ whose direct sum with $O_{+}$ is trivial, and form $(\alpha^{\prime}:\pi^{\ast}O^{\prime}_{+}\to\pi^{\ast}O^{\prime}_{-})=(\alpha\oplus 1:\pi^{\ast}(O_{+}\oplus O^{\prime})\to\pi^{\ast}(O_{-}\oplus O^{\prime}))$. Since $\alpha^{\ast}\alpha$ is multiplication by $|\xi|^{2}+|s(x)|^{2}$, $\alpha$, hence $\alpha^{\prime}$, is a bundle isomorphism outside the compact set $s^{-1}(0)$. Under a suitable trivialization of $O^{\prime}_{\pm}$, $\alpha^{\prime}$ is the identity map near $\partial Y$. We can also modify $\alpha$ so that it is constant along $\{(x,\nu\xi):\nu\in\mathbb{R}_{+}\}$ over the entire $Y$ by [4] page 492-493.
Let $H_{\pm}=L^{2}(Y,O^{\prime}_{\pm})$. Note that $H_{\pm}$ carry the structure of $\mathbb{C}(Y)$-modules via the natural multiplication by continuous $\mathbb{C}$-valued functions on the base, which can be understood as *-representations $\phi_{\pm}:\mathbb{C}(Y)\to L(H_{\pm})$ to the algebras of bounded operators on the Hilbert spaces. Let $T_{s}:H_{+}\to H_{-}$ be the order 00 pseudo-differential operator obtained by continuously extending the quantization of $\alpha^{\prime}$ from compactly supported smooth sections of $O^{\prime}_{\pm}$ to $H_{\pm}$. The datum $(H_{\pm},\phi_{\pm},T_{s})$ defines an element in Kasparov’s K-homology group $K_{0}(Y)$, which we will denote as $\mathcal{O}^{vir}_{s}$ and refer to as the K-theoretic virtual fundamental class.
In particular, if $s$ is transverse to the zero section, so that $M=s^{-1}(0)$ is a stably almost complex submanifold, such extension construction allows us to define the K-theoretic fundamental class $[M]\in K_{0}(\Sigma^{N}M^{+})=K_{0}(M)$.
@@
Let $i:M\to Y$ be the inclusion map. Then $i_{*}[M]=\mathcal{O}^{vir}_{s}$.
Given a virtual bundle $V=V_{+}\ominus V_{-}\in K^{0}(Y)$, we define the pairing $\mathcal{O}^{vir}_{s},V\right>$ as the difference $n_{+}-n_{-}$ of Fredholm indices
$\begin{array}[]{ll}n_{+}&=\operatorname{ind}(T_{s}\otimes 1_{+}:H_{+}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{+})\to H_{-}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{+}));\\ n_{-}&=\operatorname{ind}(T_{s}\otimes 1_{-}:H_{+}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{-})\to H_{-}\otimes_{\mathbb{C}(Y)}\mathbb{C}(V_{-})),\end{array}$
where $\mathbb{C}(V)$ is the space of complex-valued continuous sections of the bundle $V$, and has the structure of a finitely-generated $\mathbb{C}(Y)$-module.
@@
The pairing $\mathcal{O}^{vir}_{s},V\right>$ equals the index of the virtual bundle
$O\otimes\pi^{\ast}V\in K^{0}(T^{\ast}Y),$
@@
This theorem shows that our definition agrees with that in [2].
@@ 1.2. Results
In this paper, we defined a $K$-theoretic virtual fundamental class $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ of an almost complex global Kuranishi chart $(\mathcal{Y,E},\sigma)$. The $K$-theoretic virtual fundamental class is an element in $K_{0}(|I\mathcal{Y}|)$, the (complex coefficient) K-homology of the coarse space $|I\mathcal{Y}|$ of the inertia orbifold of $\mathcal{Y}$. We also define the pairing of $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ and elements in the (orbifold) K-group of $\mathcal{Y}$.
Sector-wise, $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ is constructed from $(|\mathcal{Y}^{(g)}|,str_{(g)}\mathcal{E}|_{\mathcal{Y}^{(g)}},str_{(g)}\circ\sigma|_{\mathcal{Y}^{(g)}})$ as in the no-isotropy case.
@@
  • The cap product of $(T:H_{+}\to H_{-})\in K_{0}(|I\mathcal{Y}|)$ and $(\alpha:V_{+}\to V_{-})\in K^{0}(|I\mathcal{Y}|)$ is
@@
    Here in the representatives, $H_{\pm}$ are Hilbert spaces and $V_{\pm}$ are Hermitian bundles, $C(V_{\pm})$ are continuous sections of $V_{\pm}$ and $C(\alpha):C(V_{+})\to C(V_{-})$ the induced map.
  • The map $p_{\ast}$ is an extension of the Fredholm index on $K$-homology classes represented by elliptic operators. The precise definition is in Sections 3.3-3.4 via inclusion-exclusion and Kawasaki’s index theorem [11].
@@
(Orbifold version) If $\sigma$ is transverse to the zero section, so that $\mathcal{M}=\sigma^{-1}(0)$ is a stably almost complex sub-orbifold, such extension construction allows us to define the K-theoretic fundamental class $[\mathcal{M}]\in\oplus_{(g)}K_{0}(\Sigma^{N}|\mathcal{M}^{(g)}|^{+})=K_{0}(\mathcal{M})$. Then
$i_{*}[\mathcal{M}]=\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}.$
@@
(Orbifold version) Given a virtual bundle $\mathcal{V}\in K^{0}(\mathcal{Y})$, the pairing of $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}$ with $\mathcal{V}$ equals the invariants defined by $\mathcal{V}$ in [2].
@@ 2. Orbifold K-theory
Let $\mathcal{Y}=Y/G$ be a compact orbifold (with boundary), written as a compact manifold $Y$ quotient by a compact Lie group $G$ ([13] Corollary 1.3). We use cursive letters for objects on orbifolds and printed-style letters for $G$-equivariant and $G$-invariant objects on manifolds. We use $\varepsilon$ for the ‘grading’ operator on various $\mathbb{Z}_{2}$-graded objects, i.e., the operator that acts as $(-1)^{a}$ on the subspace with grading $a$. We use $\mathbb{C}(\mathcal{Y})$ for the $\mathbb{C}^{\ast}$-algebra of complex-valued continuous functions (under the $\sup$-norm) on $\mathcal{Y}$.
@@
We recall the relative K-group $K^{0}(\mathcal{Y},\partial\mathcal{Y})$ of the orbifold $\mathcal{Y}=Y/G$.
@@
For an equivariant bundle map $\alpha:E_{+}\to E_{-}$ between complex $G$-bundles over $Y$, we define the support of $\alpha$ as the (closed) set of points in $Y$ over which $\alpha$ is not an isomorphism of representations of the isotropy groups. The set $\mathcal{C}_{G}(Y,\partial{Y})$ of all bundle map $\alpha$’s supported off $\partial Y$ is a semi-group under direct sum (defined as $\alpha^{0}\oplus\alpha^{1}:E^{0}_{+}\oplus E^{1}_{+}\to E^{0}_{-}\oplus E^{1}_{-}$ for $\alpha^{i}:E^{i}_{+}\to E^{i}_{-}$). The zero element is the identity map between 0-dimensional trivial bundles over $Y$.
@@
An element $\alpha:E_{+}\to E_{-}$ in $\mathcal{C}_{G}(Y,\partial{Y})$ is degenerated if it has empty support. Two elements $\alpha^{0}:E^{0}_{+}\to E^{0}_{-}$ and $\alpha^{1}:E^{1}_{+}\to E^{1}_{-}$ in $\mathcal{C}_{G}(Y)$ are homotopic if there is an element $\tilde{\alpha}:\tilde{E}_{+}\to\tilde{E}_{-}$ in $C_{G}(Y\times[0,1],(\partial Y)\times[0,1])$ that restricts to $\alpha^{0}$ and $\alpha^{1}$ respectively at $Y\times\{0\}$ and $Y\times\{1\}$. Remark that the pair $(Y\times[0,1],(\partial Y)\times[0,1])$ equivariantly retracts to $(Y,\partial Y)$, thus we can take $\tilde{E}_{\pm}$ be the pull-back of a certain bundle over $(Y,\partial Y)$, i.e., independent of $t\in[0,1]$.
@@
We define an equivalence relation on $\mathcal{C}_{G}(Y,\partial{Y})$ generated by identifying degenerate elements with 00 and identifying homotopic elements. We define the $G$-equivariant $K$-group $K_{G}^{0}(Y,\partial{Y})$ of $Y$ as the set of equivalence classes of this equivalence relation under direct sum.
@@
In literature, sometimes we define the K-group $K_{G}^{0}(Y,\partial Y)$ as follows. We define the absolute $K$-groups $K^{0}_{G}$ of a space as the Grothendieck group associated to the semi-group (under direct sum) of isomorphism classes of finite-dimensional complex $G$-bundles over that space. A map between spaces induces a pull-back of bundles, which is a homomorphism between corresponding absolute $K$-groups. We define the reduced $K$-group $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$ as the kernel of the pull-back of $K^{0}_{G}(Y/\partial Y)\to K^{0}_{G}(pt)$ along the inclusion $pt\to Y/\partial Y$ of the base point, and define $K_{G}^{0}(Y,\partial Y)=\tilde{K}_{G}^{0}(Y/\partial Y,pt)$.
@@
Note that by definition, each element in $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$ is represented by some element with $E_{+}|_{U}=E_{-}|_{U}=U\times_{G}V$, for some representation $V$ of $G$, in a $G$-invariant neighborhood $U$ of $\partial Y$. This definition coincides with our first definition using compactly supported bundle homomorphisms ([14] Proposition 3.1, see also [4] page 490). An isomorphism is explicitly constructed as follows. Given $E_{+}\ominus E_{-}\in\tilde{K}_{G}^{0}(Y/\partial Y,pt)$, let $\alpha:E_{+}\to E_{-}$ be an arbitrary map that is the identity map in $U$. The pair $(E_{\pm},\alpha)$ represents an element in $K^{0}_{G}(Y,\partial Y)$. Conversely, given a bundle map $\alpha:E_{+}\to E_{-}$, we pick a bundle $E$ over $Y$, and a $G$-invariant neighborhood $U$ of $\partial Y$, such that $(E_{+}\oplus E)|_{U}$ takes form $U\times_{G}V$. Then $\alpha\oplus id:E_{+}\oplus E\to E_{-}\oplus E$ represents an element in $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$.
@@
For notation simplicity, we define $K^{0}_{G}(Y,\partial{Y})$ as pairs $(E,\alpha)$ of a $\mathbb{Z}_{2}$-graded Hermitian $G$-bundle $E$ and a self-adjoint degree $1$ compactly supported bundle automorphism $\alpha:E\to E$, modulo degenerate elements and homotopy. Indeed, $(E_{+},E_{-},\alpha_{+})$ in the first definition corresponds to $E=E_{+}\oplus E_{-}$ and the map
@@
Here we identify $(E_{\pm})^{\ast}$ with $E_{\pm}$ by picking a Hermitian metric on $E_{\pm}$. Hermitian metrics on $E_{\pm}$ define homotopic representatives of the $K$-class, thus the same elements in the $K$-group.
The orbifold K group is defined as $K^{0}(Y/G,\partial{Y}/G)=K_{G}^{0}(Y,\partial{Y})$.
2.2. Twisted sectors, inertia orbifold, and fake K-group
Given a quotient orbifold $\mathcal{Y}=Y/G$, we define the inertia orbifold

### v2

# K-theoretic Virtual Fundamental Cycle of a Global Kuranishi Chart

@@ Abstract.
In this paper, we define the virtual fundamental cycle of a global Kuranishi chart as an element in the (analytic) orbispace K-homology of the virtual orbifold and verify that it defines the same invariants as those in [2].
@@ 1. Introduction
The motivation of this paper is to define permutation-equivariant quantum K-theory [12][6][7] for general symplectic target spaces. The essential ingredient in this definition is the $K$-theoretical virtual fundamental cycle of moduli spaces $\mathcal{M}$ of $J$-holomorphic curves. There are various theories characterizing these moduli spaces. We use almost complex global Kuranishi charts [1][2] in this paper.
@@
In this paper, we define $K$-theoretical virtual fundamental cycles as elements in the $K$-homology of certain spaces, instead of the $K$-cohomology of cotangent spaces as in [1][2].
@@ 1.1. The case of manifolds
Let $Y$ be a compact almost complex manifold (possibly with boundary), and let $s$ be a (continuous, not necessarily smooth) section of a complex vector bundle $E$ over $Y$ whose zero locus $M=s^{-1}(0)$ lies in the interior of $Y$. Pick Hermitian metrics on $T^{\ast}Y$ and $E$.
Let $\pi:T^{\ast}Y\to Y$ and define
$\begin{array}[]{llll}S_{+}&=\wedge^{\text{even}}T^{\ast}Y,&S_{-}&=\wedge^{\text{odd}}T^{\ast}Y,\\ E_{+}&=\wedge^{\text{even}}E,&E_{-}&=\wedge^{\text{odd}}E.\end{array}$
@@
Pick a vector bundle $O^{\prime}$ whose direct sum with $O_{+}$ is trivial, and form $(\alpha^{\prime}:\pi^{\ast}O^{\prime}_{+}\to\pi^{\ast}O^{\prime}_{-})=(\alpha\oplus 1:\pi^{\ast}(O_{+}\oplus O^{\prime})\to\pi^{\ast}(O_{-}\oplus O^{\prime}))$. Since $\alpha^{\ast}\alpha$ is multiplication by $|\xi|^{2}+|s(x)|^{2}$, $\alpha$, and hence $\alpha^{\prime}$, is a bundle isomorphism outside the compact set $s^{-1}(0)$. Under a suitable trivialization of $O^{\prime}_{\pm}$, $\alpha^{\prime}$ is the identity map near $\partial Y$. We can also modify $\alpha$ so that it is constant along $\{(x,\nu\xi):\nu\in\mathbb{R}_{+}\}$ over the entire $Y$ by [4], pages 492–493.
Let $H_{\pm}=L^{2}(Y,O^{\prime}_{\pm})$. Note that $H_{\pm}$ carry the structure of $\mathbb{C}(Y)$-modules via natural multiplication by continuous $\mathbb{C}$-valued functions on the base, which can be understood as *-representations $\phi_{\pm}:\mathbb{C}(Y)\to L(H_{\pm})$ into the algebras of bounded operators on the Hilbert spaces. Let $T_{s}:H_{+}\to H_{-}$ be the order 0 pseudo-differential operator obtained by continuously extending the quantization of $\alpha^{\prime}$ from compactly supported smooth sections of $O^{\prime}_{\pm}$ to $H_{\pm}$. Let $i_{!}:\mathbb{C}(M)\to\mathbb{C}(Y)$ be an arbitrary continuous linear operator, such that when restricted to each point, it is a continuous extension of the corresponding function. The datum $(H_{\pm},\phi_{\pm}\circ i_{!},T_{s})$ defines an element in Kasparov’s K-homology group $K_{0}(M)$, which is independent of $i_{!}$ and the global chart $(Y,E,s)$. We denote this element as $\mathcal{O}^{vir}$ and refer to it as the K-theoretic virtual fundamental class.
In particular, if $s$ is transverse to the zero section so that $M=s^{-1}(0)$ is a stably almost complex submanifold, such an extension construction allows us to define the K-theoretic fundamental class $[M]\in K_{0}(\Sigma^{N}M^{+})=K_{0}(M)$.
@@
$[M]=\mathcal{O}^{vir}$.
Given a virtual bundle $V=V_{+}\ominus V_{-}\in K^{0}(M)$, we define the pairing $\mathcal{O}^{vir},V\right>$ as the difference $n_{+}-n_{-}$ of Fredholm indices:
$\begin{array}[]{ll}n_{+}&=\operatorname{ind}(T_{s}\otimes 1_{+}:H_{+}\otimes_{\mathbb{C}(M)}\mathbb{C}(V_{+})\to H_{-}\otimes_{\mathbb{C}(M)}\mathbb{C}(V_{+}));\\ n_{-}&=\operatorname{ind}(T_{s}\otimes 1_{-}:H_{+}\otimes_{\mathbb{C}(M)}\mathbb{C}(V_{-})\to H_{-}\otimes_{\mathbb{C}(M)}\mathbb{C}(V_{-})),\end{array}$
where $\mathbb{C}(V)$ is the space of complex-valued continuous sections of the bundle $V$ and has the structure of a finitely generated $\mathbb{C}(M)$-module.
In particular, if $V=i^{\ast}\tilde{V}$ for $i:M\to Y$ the inclusion, and $\tilde{V}\in K^{0}(Y)$, then we have the following theorem.
@@
The pairing $\mathcal{O}^{vir},V\right>$ equals the index of the virtual bundle
$O\otimes\pi^{\ast}\tilde{V}\in K^{0}(T^{\ast}Y),$
@@
This theorem shows that our definition of invariants agrees with that in [2].
@@ 1.2. Results
In this paper, we define a $K$-theoretic virtual fundamental class $\mathcal{O}^{vir}$ of an almost-complex global Kuranishi chart $(\mathcal{Y},\mathcal{E},\sigma)$. The $K$-theoretic virtual fundamental class is an element in $K_{0}(|I\mathcal{M}|)$, the (complex-coefficient) $K$-homology of the coarse space $|I\mathcal{M}|$ of the inertia orbispace of $\mathcal{M}$. We also define the pairing of $\mathcal{O}^{vir}$ with elements in the (orbispace) $K$-group of $\mathcal{M}$.
Sector-wise, $\mathcal{O}^{vir}$ is constructed from $(|\mathcal{Y}^{(g)}|,str_{(g)}\mathcal{E}|_{\mathcal{Y}^{(g)}},str_{(g)}\circ\sigma|_{\mathcal{Y}^{(g)}})$ as in the no-isotropy case.
@@
  • The cap product of $(T:H_{+}\to H_{-})\in K_{0}(|I\mathcal{M}|)$ and $(\alpha:V_{+}\to V_{-})\in K^{0}(|I\mathcal{M}|)$ is
@@
    Here, in these representatives, $H_{\pm}$ are Hilbert spaces, $V_{\pm}$ are Hermitian bundles, $C(V_{\pm})$ are continuous sections of $V_{\pm}$, and $C(\alpha):C(V_{+})\to C(V_{-})$ is the induced map.
  • For orbifolds, the map $p_{\ast}$ is an extension of the Fredholm index on $K$-homology classes represented by elliptic operators. The precise definition is given in Section 3.5 via inclusion-exclusion and Kawasaki’s index theorem [11]. Its definition is extended to virtual orbifolds in Section 4.2.
@@
(Orbifold version) If $\sigma$ is transverse to the zero section, so that $\mathcal{M}=\sigma^{-1}(0)$ is a stably almost complex sub-orbifold, such an extension construction allows us to define the $K$-theoretic fundamental class $[\mathcal{M}]\in\oplus_{(g)}K_{0}(\Sigma^{N}|\mathcal{M}^{(g)}|^{+})=K_{0}(\mathcal{M})$. Then
$[\mathcal{M}]=\mathcal{O}^{vir}.$
@@
(Orbifold version) Let $i:\mathcal{M}=\sigma^{-1}(0)\to\mathcal{Y}$ be the inclusion. Given a virtual bundle $\mathcal{V}\in K^{0}(\mathcal{Y})$, the pairing of $\mathcal{O}^{vir}$ with $i^{\ast}\mathcal{V}$ equals the invariants defined by $\mathcal{V}$ in [2].
@@ 2. Orbifold K-theory
Let $\mathcal{Y}=Y/G$ be a compact orbifold with boundary (unless otherwise mentioned), which is the quotient of a compact manifold $Y$ by a compact Lie group $G$ ([13] Corollary 1.3). We use cursive letters for objects on orbifolds and printed-style letters for $G$-equivariant and $G$-invariant objects on manifolds. We use $\varepsilon$ for the grading operator on various $\mathbb{Z}_{2}$-graded objects, i.e., the operator that acts as $(-1)^{a}$ on the subspace of grading $a$. We use $\mathbb{C}(\mathcal{Y})$ for the $\mathbb{C}^{\ast}$-algebra of complex-valued continuous functions (under the $\sup$-norm) on $\mathcal{Y}$.
@@
We recall the relative K-group $K^{0}(\mathcal{Y},\partial\mathcal{Y})$ of the orbispace $\mathcal{Y}=Y/G$. Note that in this paper the orbispace is either an orbifold (the base space of the global Kuranishi chart) or a virtual orbifold (the zero locus of the section in the global Kuranishi chart). In the latter case, $\partial\mathcal{Y}=\emptyset$.
@@
For an equivariant bundle map $\alpha:E_{+}\to E_{-}$ between complex $G$-bundles over $Y$, we define the support of $\alpha$ as the (closed) set of points in $Y$ where $\alpha$ is not an isomorphism of the representations of the isotropy groups. The set $\mathcal{C}_{G}(Y,\partial{Y})$ of all bundle maps $\alpha$ supported away from $\partial Y$ is a semigroup under direct sum (defined as $\alpha^{0}\oplus\alpha^{1}:E^{0}_{+}\oplus E^{1}_{+}\to E^{0}_{-}\oplus E^{1}_{-}$ for $\alpha^{i}:E^{i}_{+}\to E^{i}_{-}$). The zero element is the identity map between 0-dimensional trivial bundles over $Y$.
@@
An element $\alpha:E_{+}\to E_{-}$ in $\mathcal{C}_{G}(Y,\partial{Y})$ is called degenerate if it has empty support. Two elements $\alpha^{0}:E^{0}_{+}\to E^{0}_{-}$ and $\alpha^{1}:E^{1}_{+}\to E^{1}_{-}$ in $\mathcal{C}_{G}(Y,\partial{Y})$ are homotopic if there exists an element $\tilde{\alpha}:\tilde{E}_{+}\to\tilde{E}_{-}$ in $C_{G}(Y\times[0,1],(\partial Y)\times[0,1])$ that restricts to $\alpha^{0}$ and $\alpha^{1}$ at $Y\times\{0\}$ and $Y\times\{1\}$, respectively. Note that the pair $(Y\times[0,1],(\partial Y)\times[0,1])$ equivariantly retracts to $(Y,\partial Y)$; thus we can take $\tilde{E}_{\pm}$ to be the pull-back of a certain bundle over $(Y,\partial Y)$, i.e., independent of $t\in[0,1]$.
@@
We define an equivalence relation on $\mathcal{C}_{G}(Y,\partial{Y})$ generated by identifying degenerate elements with 00 and by identifying homotopic elements. We define the $G$-equivariant $K$-group $K_{G}^{0}(Y,\partial{Y})$ of $Y$ as the set of equivalence classes of this relation under direct sum.
@@
In the literature, the $K$-group $K_{G}^{0}(Y,\partial Y)$ is sometimes defined as follows. We define the absolute $K$-groups $K^{0}_{G}$ of a space as the Grothendieck group associated to the semigroup (under direct sum) of isomorphism classes of finite-dimensional complex $G$-bundles over that space. A map between spaces induces a pull-back of bundles, which is a homomorphism between the corresponding absolute $K$-groups. We define the reduced $K$-group $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$ as the kernel of the pull-back map $K^{0}_{G}(Y/\partial Y)\to K^{0}_{G}(pt)$ induced by the inclusion $pt\to Y/\partial Y$ of the base point, and define $K_{G}^{0}(Y,\partial Y)=\tilde{K}_{G}^{0}(Y/\partial Y,pt)$.
@@
Note that, by definition, each element in $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$ is represented by an element for which $E_{+}|_{U}=E_{-}|_{U}=U\times_{G}V$, for some representation $V$ of $G$, on a $G$-invariant neighborhood $U$ of $\partial Y$. This definition coincides with our first definition using compactly supported bundle homomorphisms ([14] Proposition 3.1, see also [4] page 490). An isomorphism is explicitly constructed as follows. Given $E_{+}\ominus E_{-}\in\tilde{K}_{G}^{0}(Y/\partial Y,pt)$, let $\alpha:E_{+}\to E_{-}$ be an arbitrary map that restricts to the identity on $U$. The pair $(E_{\pm},\alpha)$ represents an element in $K^{0}_{G}(Y,\partial Y)$. Conversely, given a bundle map $\alpha:E_{+}\to E_{-}$, we pick a bundle $E$ over $Y$, and a $G$-invariant neighborhood $U$ of $\partial Y$, such that $(E_{+}\oplus E)|_{U}$ is of the form $U\times_{G}V$. Then $\alpha\oplus id:E_{+}\oplus E\to E_{-}\oplus E$ represents an element in $\tilde{K}_{G}^{0}(Y/\partial Y,pt)$.
@@
For simplicity, we define $K^{0}_{G}(Y,\partial{Y})$ as pairs $(E,\alpha)$ where $E$ is a $\mathbb{Z}_{2}$-graded Hermitian $G$-bundle and $\alpha:E\to E$ is a self-adjoint, degree $1$, compactly supported bundle automorphism, modulo degenerate elements and homotopy. Indeed, $(E_{+},E_{-},\alpha_{+})$ in the first definition corresponds to $E=E_{+}\oplus E_{-}$ and the map
@@
Here, we identify $(E_{\pm})^{\ast}$ with $E_{\pm}$ by choosing a Hermitian metric on $E_{\pm}$. Different choices of Hermitian metrics on $E_{\pm}$ yield homotopic representatives of the $K$-class, and hence the same element in the $K$-group.
The orbispace K group is defined as $K^{0}(Y/G,\partial{Y}/G)=K_{G}^{0}(Y,\partial{Y})$.
2.2. Twisted sectors, inertia orbispace, and fake K-group
Given a quotient orbispace $\mathcal{Y}=Y/G$, we define the inertia orbispace